A 5.0 L ballon has a pressure of 1.0 atm at sea level, where the temperature is 28°C. The balloon was let loose and ascended to a height in the temperature where the temperature is 11°C. What will be the volume of the balloon if the atmospheric pressure at the new altitude is 0.65 atm?

1
Expert's answer
2022-02-25T08:05:01-0500

P1V1/T1=P2V2/T2

V2=V1 (T2P1/T1P2)

V1=(273+11)(1)/(273+28)(0.65)

=1.45V1